what is the longest line you can draw a ruler on a piece of paper 20cm by 30cm? can you please send me the question as soon as possible please!!!!!!THANK YOU
Answer by khwang(438)

The longest lien (segement ) on this paper of rectangular shape is
the diagonal D which length is [since D^2 = 20^ + 30^2, why?]
sqrt(20^ + 30^2) = sqrt(1300) = 10sqrt(13)